OpenStudy (anonymous):

how can I prove that d/dx (1/x) = -1/x^2. ??? how can I prove that d/dx (1/x) = -1/x^2. ??? @MIT 18.01 Single …

OpenStudy (anonymous):

1/x <=> x^-1 d?dx x^-1 = -1x^-2, = -1/x^2 qed

OpenStudy (anonymous):

I will prove that using the definition in the lecture. Question: Prove that \[d/dx (1/x) = -1/ x^{2}\]Answer: \[f(x)=1/x\]\[d/dx(1/x)=\lim_{\Delta x \rightarrow 0} (f(x _{0}+\Delta x)-f(x _{0}))/\Delta x\](Notice that we have not been able to substitute Delta x=0 in)\[=\lim_{\Delta x \rightarrow 0}(1/(x_{0}+\Delta x)-1/x_{0})/\Delta x\]\[= \lim_{\Delta x \rightarrow 0}(x_{0}-(x_{0}+\Delta x))/((x_{0}+\Delta x).x_{0}.\Delta x)\]\[=\lim_{\Delta x \rightarrow 0}(-\Delta x)/((x_{0}+\Delta x).x_{0}.\Delta x)\](Now we cancel out Delta x)\[=\lim_{\Delta x \rightarrow 0}(-1)/((x_{0}+\Delta x).x_{0})\](Now, we can substitute Delta x=0)\[=-1/x ^{2}\]